### v1

# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

Abstract
In this work, we study the gravitating lensing Euler-Heisenberg black hole surrounded by perfect fluid dark matter. This kind of black hole solution enables us to investigate the nontrivial interplay between the dark matter effects and nonlinear electrodynamics effects (or quantum electrodynamics effects). The important observable in gravitational lensing observations are calculated and discussed, including the gravitational deflection angle of light, time delay of light, precession angle of massive orbit’s bound orbits, and black hole shadow radius. The results indicates that the Euler-Heisenberg black hole with a negative perfect fluid dark matter parameter ($\lambda_{\text{DM}}<0$) could greatly intensify the gravitational deflection angle of light, time delay of light and precession angle of massive object’s bound orbit, while the nonlinear electrodynamics effects do not have large influences on these gravitational lensing observables.
Keywords: Euler Heisenberg Black Hole; Gravitational Lensing; Perfect Fluid Dark Matter;
Nonlinear Electrodynamics
1 Introduction
In the past few decades, modern astrophysical observations reveal that most of our universe is dominated by dark matter and dark energy. Particularly, the cosmic microwave background (CMB) observations suggested that $26.8\%$ of our universe is consist of dark patter, $68.3\%$ of our universe is composed of dark energy, while only $4.9\%$ of our universe is composed of baryonic matter [2, 3, 4, 5]. Theoretically, the unknown dark matter are composed of unknown particles predicted in theories beyond the Standard Model, such as axions, sterile neutrinos and weakly interactive massive particles(WIMPs). In astronomy, the dark matter have extremely strong influences on the structure and formation of galaxy clusters, and it can also dominantly affects the star motions in galaxies and galaxy clusters, producing observational flat rotational curves in spiral galaxies [6, 7, 8, 9]. Notably, recent studies also suggested that galactic dark matter halo have non-negligible influences on the gravitational lensing and shadow images of supermassive black holes [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
The equation of state for dark matter assembled in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted large interests [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trial effects and characters for dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and the effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties are extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
Inspired by the simplicity and success of PFDM model in black hole physics and astronomy, we choose a novel black hole solution surrounded by PFDM — the Euler-Heisenberg black hole. The Euler-Heisenberg black hole, which is obtained from the Euler-Heisenberg effective action [53, 54, 55, 56], stimulated numerous studies in past few years, for its nontrivial realization on the coupling between nonlinear electrodynamics and gravity systems, especially with the quantum loop-corrections in quantum electrodynamics (QED) [57]. With these desired properties, this kind of black hole is often used to explore the non-linear electrodynamics in black hole physics, resulting in a number of studies on thermodynamic transitions, thermodynamic topology, black hole shadow, gravitational lensing, quasi-normal modes for Euler-Heisenberg black hole in recent years [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 82, 78, 79, 80, 81, 86, 83, 84, 88, 85, 87]. The nonlinear electrodynamics effects (or QED effects) on gravitation are extremely important issues in high energy scale, and they may have potential influences on the evolution of our universe (especially at the beginning stage), such as the inflationary theory in Big Bang, the dynamic evolution of dark matter or dark energy, and the cosmological accelerating expansion of universe. Therefore, it is interesting to study the interplay between dark matter and non-linear electrodynamics in gravitational systems. The Euler-Heisenberg black hole surrounded by PFDM could provide us a simply way to investigate their influences on charged black holes. Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [89, 90, 91, 92], which enable us to give a comprehensive exploration on the dark matter effects and non-linear electrodynamics effects (or QED effects) in black hole physics.
In this work, we carry out an investigation on the gravitational lensing of Eular-Heisenberg black hole surrounded by PFDM. The gravitational lensing observables provide a avenue to reveal the features of gravitational field and intrinsic properties of black holes through high precision observational data [93]. And an exploration of gravitational lensing observables for Eular-Heisenberg black hole surround by PFDM could enrich our knowledge on the gravitational field for charged black hole systems, especially the influences coming from the dark matter effects and nonlinear electrodynamics effects (or QED effects). Concretely, in the present work, we calculate the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits for Euler-Heisenberg black hole surrounded by PFDM. There are significantly important observables in the gravitational lensing observations, which have been frequently used to test various aspects of black hole physics and gravitational theories in both strong gravitational field limit and weak gravitational field limit. Furthermore, stimulated by the rapid development of black hole shadow since the capture of black hole images by Event Horizon Telescope (EHT) [94, 95, 96], the black hole critical shadow radius for Euler-Heisenberg black hole in the presence of PFDM is also discussed.
This rest of the present work is organized in the following way. Section 2 briefly introduces the spacetime metric generated by Euler-Heisenberg black hole surrounded by PFDM. In section 3, the theoretical approach we employed in calculating the gravitational lensing observables is described in detailed. In section 4, results and discussions on the gravitational lensing observables (gravitational deflection angle of light, time delay of light, trajectory and precession angle of massive object’s bound orbit, black hole shadow critical radius) for Euler-Heisenberg black holes surrounded by PFDM are presented. Summary and perspectives are given in section 5. Throughout this work, the natural unit $G=c=1$ is adopted.
2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
The Euler-Heisenberg black hole becomes attractive in recent year for its ability to include the nonlinear electrodynamics effects (or QED effects). This kind of black hole can be solved from the gravitational field equation and the Euler-Heisenberg effective action [53], and it is one of the most important realization on the coupling between nonlinear electrodynamics and charged black hole systems.
Recently, several analytical solution for Euler-Heisenberg black hole in the presence of dark matter medium have been constructed [90, 89, 91, 92]. Particularly, S.-J. Ma et al. derived the effective spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM, through solving the gravitational field equation coupled with nonlinear electromagnetic field (with QED effects included) and dark matter field (assumed to be PFDM) [89]. The Lagrangian of a gravitational system coupled with nonlinear electromagnetic field and dark matter field can be written as [89, 91]
$S=\frac{1}{4\pi}\int\sqrt{-g}d^{4}x\cdot\bigg{[}\frac{R}{4}+\mathcal{L}_{\text{EM}}+\mathcal{L}_{\text{PFDM}}\bigg{]}$ (1)
and the nonlinear electromagnetic field’s Lagrangian density is given by
$\mathcal{L}_{\text{EM}}=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2}$ (2)
where $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is electromagnetic strength tensor and $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $*F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The above Lagrangian leads to the gravitational field equation
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu}$ (3)
with the energy-momentum tensor contributed from nonlinear electromagnetic field and dark matter field. To derive a spherically symmetric analytical solution for black hole, the energy-momentum tensor for PFDM and nonlinear electromagnetic field are restricted to the following explicit form in the spherical coordinates [89]
$T^{\mu}_{\nu}$ $=$ $(T^{t}_{t},\ T^{r}_{r},\ T^{\theta}_{\theta},\ T^{\phi}_{\phi})$ (4)
$=$ $\frac{1}{4\pi}\bigg{(}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}}\bigg{)}\ \ \text{Electromagnetic Field}$
$=$ $\bigg{(}\frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ \frac{\lambda_{\text{DM}}}{8\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}},\ -\frac{\lambda_{\text{DM}}}{16\pi r^{3}}\bigg{)}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{Dark Matter Field}$
where $M$ is the black hole mass, $Q$ is the black hole electric charge, $a$ is a parameter which measures the strength of non-linear electrodynamics effects (or QED effects). The parameter $\lambda_{\text{DM}}$ denotes the influences of PFDM, which is proportional to the dark matter density. Eventually, by solving the gravitational field equation in the spherically symmetric coordinates, the spacetime metric for an Euler-Heisenberg black hole surrounded by PFDM is
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$ (5)
with the metric component function $f(r)$ defined as [89]
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}$ (6)
In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric turn to be the Euler-Heisenberg black hole with non-linear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects / QED effect vanishes and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of Euler-Heisenberg black hole can be greatly influenced. It is commonly known that the classical charge black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a\neq 0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge. Therefore, in the rest of this work, it is safely to choose black hole parameters to study the gravitational lensing observables, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in the Appendix A.
3 Theoretical Methods to Obtain Gravitational Lensing Observables
This section gives descriptions on the theoretical methods used to calculate the gravitational lensing observables in our work. The treatments of particle orbits (which are geodesics) are presented in subsection 3.1. Schemes to obtained gravitational deflection angle of light and time delay of light for finite distance light source and observer are introduced in subsection 3.2 - 3.3. Finally, the formulations on the trajectories and precession angle for massive object’s bound orbits around black holes are given in subsection 3.4.
3.1 Geodesics
In the gravitational filed, massive and massless particles are moving along geodesics in 4 dimensional spacetime. Here we provide a derivation of the reduced geodesic equation, which turned out to be extremely useful in the exploration of gravitational lensing observables. For a spherically symmetric spacetime with the metric form
$d\tau^{2}=f(r)dt^{2}-\frac{1}{f(r)}dr^{2}-r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})$ (7)
the following conserved quantities can be introduced from the symmetry and Killing vectors of spacetime
$J$$\equiv$ $r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$ (8a)
$E$$\equiv$ $f(r)\frac{dt}{d\lambda}$ (8b)
$\epsilon$$\equiv$ $g_{\mu\nu}dx^{\mu}dx^{\nu}=f(r)\bigg{(}\frac{dt}{d\lambda}\bigg{)}^{2}-\frac{1}{f(r)}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}-r^{2}\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2}-r^{2}\sin^{2}\theta\bigg{(}\frac{d\phi}{d\lambda}\bigg{)}^{2}$ (8c)
Here, $\lambda$ is an affine parameter in a particle orbit, $J$ is the conserved angular momentum along the particle orbit, and $E^{2}/2$ can be viewed as the conserved energy along a particle orbit. Specifically, for spherically symmetric spacetime, we can always assume that test particles are moving in the equatorial plane $\theta=\pi/2$, and the reduced differential equation can be obtained using these conserved quantities
$\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+\frac{1}{2}f(r)\bigg{[}\frac{J^{2}}{r^{2}}+\epsilon\bigg{]}=\frac{1}{2}\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+V_{\text{eff}}(r)=\frac{1}{2}E^{2}$ (9)
Here, $V_{\text{eff}}(r)=\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}$ is the effective potential of test particles moving in spherically symmetric gravitational field, and $b\equiv|J/E|$ is the impact parameter in a particle orbit (which is also conserved along this orbit). For massless particles traveling along null geodesics, the parameter $\epsilon$ takes the value $\epsilon=0$. For massive particles traveling along timelike geodesics, the parameter $\epsilon$ takes the value $\epsilon=1$.
3.2 Gravitational Deflection Angle of Light
The gravitational deflection angle reflect the bending or distortion effects on particle orbit in a gravitational field. In the gravitational lensing observation, when the light source and observer are located at a finite distance region to the central black hole (with the radial coordinate $r=r_{\text{S}}$ and $r=r_{\text{O}}$, respectively), the gravitational deflection angle of light is defined through [97]
$\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$ (10)
The notations $\Psi_{\text{O}}$, $\Psi_{\text{S}}$ denote the angles between light ray propagation direction and the radial direction (which are measured at the observer’s position and source position), and $\Delta\phi_{\text{OS}}$ is the variation of azimuthal angle $\phi$ in the light trajectory. For a spherically symmetric spacetime described by metric expression (7), the angle between light ray propagation direction and the radial direction is uniquely determined by spacetime metric [97], via
$\sin\Psi=\frac{b\cdot\sqrt{f(r)}}{r}$ (11)
In the thin lens approximation, this definition of gravitational deflection angle for finite distance source and observer is illustrated in figure 1. The thin lens approximation assumes that the spacetime is flat everywhere except for the location of thin lens (the central black hole). For the quadrangle depicted in figure 1, the exterior angles at each point are $\Psi_{\text{S}}$, $\alpha$, $\pi-\Psi_{\text{O}}$, $\pi-\Delta\phi_{\text{OS}}$, respectively. Under the thin lens approximation, the sum of exterior angles for the quadrangle equals $2\pi$, which justify the definition of gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$. However, it should be noted that this definition given by Ishihara et al. is generally valid in gravitational lensing treatments which goes beyond the thin lens approximation, see references [97, 98, 99, 100, 101] for more discussions on this issue.
Figure 1: The gravitational deflection angle of light for finite distance light source and observer. This figure illustrate the gravitational deflection angle $\alpha\equiv\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$ in the thin lens approximation.
To calculate the gravitational deflection angle of light, it is necessary to find the variation of azimuthal angle $\phi$ in the photon orbit. In a spherically symmetric spacetime, the variation of azimuthal angle $\phi$ is derived from the reduced differential equation (9) for null geodesics
$\frac{d\phi}{dr}=\frac{d\phi}{d\lambda}\cdot\frac{d\lambda}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}$ (12)
where we have used the affine parameter $\epsilon=0$ for massless photon, the conserved energy parameter $E=f(r)\frac{dt}{d\lambda}$, conserved angular momentum $J=r^{2}\sin^{2}\theta\frac{d\phi}{d\lambda}$, and the definition of impact parameter $b\equiv|J/E|$. The plus and minus sign $\pm$ can be determined in the following way. When the photon beams travel along a scattering orbit from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as the azimuthal angle $\phi$ increases, until photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase with the increasing of azimuthal angle.
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}<0\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point$r=r_{0}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}>0\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position$r=r_{\text{O}}$}$
For the closet distance to central black hole in the photon orbit, the derivative $dr/d\lambda$ in equation (9) vanishes automatically
$\frac{dr}{d\lambda}\bigg{|}_{r=r_{0}}=0\ \ \ \Rightarrow\ \ \ \frac{1}{2}f(r_{0})\frac{J^{2}}{r_{0}^{2}}=\frac{1}{2}E^{2}\ \ \ \Rightarrow\ \ \ b^{2}=\frac{J^{2}}{E^{2}}=\frac{r_{0}^{2}}{f(r_{0})}$ (14)
Given the impact parameter $b$ for a photon orbit, the closet distance $r_{0}$ can be solved from this equation.
With the expression of angle $\Psi$ in equation (11) and the variation of azimuthal angle, the gravitational deflection angle of light can be expressed as
$\alpha$ $=$ $\Psi_{\text{O}}-\Psi_{\text{S}}-\Delta\phi_{\text{OS}}$ (15)
[…]
$=$ $\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{R}})}}{r_{\text{R}}}\bigg{)}+\arcsin\bigg{(}\frac{b\sqrt{f(r_{\text{S}})}}{r_{\text{S}}}\bigg{)}-\pi+\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}$
As shown in figure 1, the angle $\Psi$ measured at the position of light source is an obtuse angle such that $\Psi_{\text{S}}=\pi-\arcsin\big{(}\frac{b\sqrt{f(r_{\text{S}})}}{r_{\text{S}}}\big{)}$.
3.3 Time Delay of Light
In the gravitational field, the elapse of time (measured by an observer) during the light propagation can be strongly affects by gravitation, and this is the time delay effect of gravitational field. To calculate the time delay of light in the gravitational lensing observation, it is necessary to computation the variation of time coordinate $t$ (measured by an distant observer) in the photon orbit. For the massless photon traveling along null geodesics (with $\epsilon=0$), one can derive the time derivative of radial coordinate from the reduced geodesic equation in (9)
$\frac{dr}{dt}=\frac{dr}{d\lambda}\cdot\frac{d\lambda}{dt}=\pm f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}$ (16)
The plus and minus sign $\pm$ can be determined similar to the calculations of gravitational deflection angle. When the photon beams emitted from the source position $r_{\text{S}}$, the radial coordinate $r$ decreases as time passes, until the photon beams reach the closest distance $r=r_{0}$ to central supermassive black hole. After passing the turning point $r=r_{0}$, the radial coordinate starts to increase as time advances.
$\frac{dt}{dr}=-\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}<0\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
$\frac{dt}{dr}=\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}>0\ \ \ \ \ \ \ \text{from tuning point $r=r_{0}$ to observer position $r=r_{\text{O}}$}$
In the gravitational lensing observation, when the light source and observer are located at $r=r_{\text{S}}$ and $r=r_{\text{O}}$, the time delay of light during the propagation process can be expressed as
[…]
$=$ $-\int_{r_{\text{S}}}^{r_{0}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}-T_{0}$
$=$ $\int_{r_{0}}^{r_{\text{S}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}+\int_{r_{0}}^{r_{\text{O}}}\frac{dr}{f(r)\sqrt{1-\frac{b^{2}f(r)}{r^{2}}}}-\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}-\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$
with $T_{0}=\sqrt{r_{\text{S}}^{2}-r_{0}^{2}}+\sqrt{r_{\text{O}}^{2}-r_{0}^{2}}$ to be the time period during the light propagation process without the gravitational field. In the expression (18), it is obvious that the time delay $\Delta T$ increases monotonically as the coordinates of light source and observer $r_{\text{S}}$, $r_{\text{O}}$ increase. In the integration process, the turning point $r=r_{0}$ is solved through equation (14).
Furthermore, it should be addressed that the time delay of light calculated using expression (18) is the time delay measured by the conventional static time coordinate $t$ in spacetime metric expression (7). For an asymptotically flat spacetime, the static time coordinate corresponds to the time measured by a very distant observer. On the other hand, if one want to calculate the time delay measured by the proper time of an finite distance static observer, a redshift factor should be taken into account
$\Delta T_{\text{observer's proper time unit}}=\frac{\Delta T}{\text{redshift factor}}=\frac{\Delta T}{\sqrt{f(r_{\text{O}})}}$ (19)
with $r_{\text{O}}$ to be the position of finite distance static observer.
3.4 Massive Object Orbits and Precession
Besides the gravitational deflection discussed in subsection 3.2 (which is relevant to a scattering orbit for massless photons), the massive object’s orbits are also greatly influenced by gravitational field, both for bound orbits and scattering orbits. Notably, massive object’s bound orbits may exhibit precession. In the gravitational lensing observations, the precession of bound orbits is frequently used to test the gravitational field around the supermassive black hole in galaxy center (such as the observation of Sgr*A in Milky Way Galaxy).
For a spherically symmetric black hole spacetime, we can always restrict the motion of massive object’s in the equatorial plane. To calculate the massive object’s orbit in the equatorial plane, it is necessary to resort to the reduced geodesic equation (9) and the definition of conserved momentum
$\bigg{(}\frac{dr}{d\tau}\bigg{)}^{2}$ $=$ $E^{2}-V_{\text{eff}}(r)=E^{2}-f(r)\bigg{[}\frac{J^{2}}{r^{2}}+1\bigg{]}$ (20a)
$\frac{d\phi}{d\tau}$ $=$ $\frac{J}{r^{2}}$ (20b)
where we have used the proper time as the affine parameter for massive objects traveling along timelike geodesics. The parameter $\epsilon$ takes the value $\epsilon=1$ for timelike geodesics. Substitute the expression (20b) into equation (20a) and cancel the affine parameter $\tau$, the geodesic equation can be further reduced to
$\bigg{(}\frac{dr}{d\phi}\frac{J}{r^{2}}\bigg{)}^{2}=E^{2}-V_{\text{eff}}(r)=E^{2}-f(r)\bigg{[}\frac{J^{2}}{r^{2}}+1\bigg{]}$ (21)
$\Rightarrow$ $\bigg{(}\frac{du}{d\phi}\bigg{)}^{2}=\frac{1}{b^{2}}-F(u)\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]}$
where we have used the new variable $u=1/r$ in the geodesic equation. Differentiating each side of this equation with respect to $\phi$, a second order differential equation for massive object’s orbit can be acquired
$\frac{d^{2}u}{d\phi^{2}}=-\frac{1}{2}\cdot\frac{dF(u)}{du}\cdot\bigg{[}u^{2}+\frac{1}{J^{2}}\bigg{]}-F(u)\cdot u$ (22)
The function $F(u)=f(1/r)$ is the spacetime metric component expressed using new variable $u=1/r$. Eventually, the massive object’s orbit is determined by numerically solving the differential equation (22).
Figure 2: A schematic plot of the precession of azimuthal angle for a massive object’s bound orbit around the central black hole.
An important feature of the massive object’s bound object is the precession of azimuthal angle $\phi$, which is illustrated in figure 2. Given a bound timelike geodesic orbit for a massive object, there are maximal and minimal position to central black hole in this orbit, denoting as $r_{\text{max}}$ and $r_{\text{min}}$. The precession of azimuthal angle $\phi$ in an orbit period can be calculated as
[…]
$=$ $2\int_{r_{\text{min}}}^{r_{max}}\frac{dr}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}-2\pi$
In the derivation of precession angle $\Delta\phi_{\text{precession}}$, we have used the derivative of of azimuthal angle with respect to radial coordinate for a timelike geodesic orbit
$\frac{d\phi}{dr}=\frac{d\phi}{d\tau}\cdot\frac{d\tau}{dr}=\pm\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}$ (24)
The plus and minus sign $\pm$ can be determined as follows. When a massive object begin to move along timelike geodesics from the maximal distance position $r_{\text{max}}$, the radial coordinate $r$ decreases as azimuthal angle $\phi$ increases, until the massive object reaches the minimal distance position $r=r_{\text{min}}$ to central supermassive black hole. After passing the minimal distance position $r=r_{\text{min}}$, the massive object’s radial coordinate starts to increase as azimuthal angle increases.
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}<0\ \ \ \text{from maximal distance $r=r_{\text{max}}$ to minimal distance $r=r_{\text{min}}$}$
$\frac{d\phi}{dr}=\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}}>0\ \ \ \ \ \text{from minimal distance $r=r_{\text{min}}$ to maximal distance $r=r_{\text{max}}$}$
Furthermore, the maximal and minimal distance position to the central black hole in a massive object’s bound orbit can be determined via
$\frac{dr}{d\phi}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}-\frac{f(r)}{J^{2}}}\bigg{|}_{r=r_{\text{max}},\ r=r_{\text{min}}}=0$ (26)
4 Results and Discussions
This section presents the results on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, including the gravitational deflection angle of light, time delay of light, the trajectory and precession angle of massive object’s bound orbits. These results are calculated numerically using expressions (15), (18) and (22). Furthermore, inspired by the rapid development of black hole shadow since the capture of black hole images by ETH, we also present a discussion on the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM.
4.1 Gravitational Deflection Angle of Light
Figure 3: The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angles are calculated numerically from expression (15) for finite distance light source and observer. This figure highlights the influences of black hole electric charge, nonlinear electrodynamics effects and dark matter effects on gravitational deflection angles. (a) The upper left panel shows the influences from black hole electric charge, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (c) The lower left panel highlights the influences form dark matter (for several PFDM parameter $\lambda_{\text{DM}}$ values) in the absence of electric charge and nonlinear electrodynamics effects ($Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the impact parameter in photon orbit, which is increased form $b=10M$ to $b=100M$. The position of light source and observer are located at a certain distance $r_{\text{O}}=r_{\text{S}}=200M$.
The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM are given in figure 3. They are calculated numerically from expression (15) for light source and observer located at finite distance region to the central black hole. In figure 3, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects / QED effects (parameterized by $a$), while the lower left and lower right panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. This figure shows the gravitational deflection in strong gravitational field cases, with the minimal impact parameter value is $b\sim 10M$, which is closer to the black hole critical shadow radius.
The upper left panel of figure 3 suggests that, for a constant PFDM parameter and non-linear electrodynamics parameter, the gravitational deflection angles are slightly reduced when black hole’s electric charge become larger. But the effect from electric charge is small such that we must add a magnified subplot into upper left panel to separate the various curves corresponding different black hole charges. The upper right panel of figure 3 suggests that the variation of nonlinear electrodynamics parameter has no obvious influences on the gravitational deflection angles, even if a magnified subplot is employed. Comparing the upper and lower panels of figure 3, we can observe that both electric charge and nonlinear electrodynamics have minor impacts on gravitational deflection angle, while the most significant influences coming from the PFDM. From the lower panels of figure 3, it can be clearly manifested that the gravitational deflection angle grow rapidly as the increasing of PFDM parameter $\lambda_{\text{DM}}$. Notably, when PFDM parameter becomes negative, the gravitational deflection angle could grow to a very large value. This can be explained from the spacetime metric in expression (6) for Euler-Heisenberg black hole surrounded by PFDM. The dark matter effects contribute to the spacetime metric function $f(r)$ via the $-\lambda_{\text{DM}}/r$ term (if we naively omit the logarithmic term of PFDM for simplicity), which is similar to the black hole’s mass term $2M/r$ (both of them are inverse proportional to the radial coordinate). This indicates that the PFDM could play a role of “effective mass” in the spacetime metric (specifically specking, because of the presence of logarithmic term, this “effective mass” should be $m_{\text{eff}}(r)=\lambda_{\text{DM}}\ln\frac{r}{|\lambda|}$, rather than $m_{\text{eff}}(r)=\lambda_{\text{DM}}$). A negative PFDM parameter ($\lambda_{\text{DM}}<0$) contributes to a positive effective mass in the spacetime metric, which explains the rapid growth of the gravitational deflection angle for negative PFDM parameter values.
4.2 Time Delay of Light
Figure 4: The gravitational time delay of light for Euler-Heisenberg black hole surrounded by PFDM. The time delays of light (which are measured by the conventional static time coordinate $t$) are calculated numerically from expression (18) for finite distance light source and observer. This figure plots the time delay results measured by conventional static time coordinate $t$, highlighting the influences coming from various black hole parameters. (a) The upper left panel shows the influences from black hole electric charge on time delay, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and dark matter parameter $\lambda_{\text{DM}}=-0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=-0.2M$. (c) The lower left panel highlights the influences form dark matter (for several PFDM parameter $\lambda_{\text{DM}}$ values) in the absence of black hole charge and nonlinear electrodynamics effects ($Q=0$ and $a=0$). (d) The lower right panel highlights the influences form dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the position of light source varied from $r_{\text{S}}=100M$ to $r_{\text{S}}=1000M$, and the vertical axis is the time delay of light $\Delta T$ in unit of black hole mass. The position of observer is chosen to be $r_{\text{O}}=200M$, and the impact parameter in photon orbit satisfies $b=10M$.
The time delays of light for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 4. This figure plot the time delay results measured by the conventional static time coordinate $t$ in spacetime metric (7). They are obtained numerically from expression (18) for light source and observer located at a finite distance region. In this figure, we plot the time delay of light varied with respect to light source position $r_{\text{S}}$. A similar plot can be carried out for time delay varied with respect to observer position $r_{\text{O}}$, while the results exhibit very similar behaviors with those in figure 4. So the plot of time delay varied with respect to observer position is not shown in this work. In figure 4, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects (or QED effects) parameterized by $a$, while the lower left and lower right panels highlight the influences form dark matter parameterized by $\lambda_{\text{DM}}$. In all four panels, the time delays of light all grow with the increasing of light source radius $r_{\text{S}}$, indicating that the time delay of light $\Delta T$ calculated using expression (18) is a monotonically function of light source radius $r_{\text{S}}$ (or observer radius $r_{\text{O}}$).
Comparing the time delay results in figure 4 and the gravitational deflection results in figure 3, it is easy to observe that the black hole parameters (including the black hole electric charge $Q$, nonlinear electrodynamics parameter $a$, PFDM parameter $\lambda_{\text{DM}}$) influence time delay of light in a similar trend as those for gravitational deflection angles. Firstly, Euler-Heisenberg black hole with a larger electric charge results in a reduction of time delay, as shown in the upper left panel of figure 4. Secondly, the nonlinear electrodynamics effects (parameterized by $a$) do not have obvious influences on time delay results, as indicated in the upper right panel of figure 4. Thirdly, the most significant impacts on time delay coming from the PFDM, which are much greater than those from electric charge and non-linear electrodynamics. From the lower panels of figure 4, the time delay of light changes rapidly as the PFDM parameter changes. Particularly, the time delay of light could grow to a very large value for a negative PFDM parameter ($\lambda_{\text{DM}}<0$), compared with those in the absence of dark matter ($\lambda_{\text{DM}}=0$). This can be explained by the “effective mass” role of PFDM in the spacetime metric, the same as we have elaborated in subsection 4.1 for gravitational deflection angle cases. A negative PFDM parameter contributes to a positive effective mass in the spacetime metric, which explains the rapid growth of the time delay for negative PFDM parameter values.
4.3 Massive Object Orbits and Precession
Figure 5: The precession angles and trajectories of massive particle orbits moving around the Euler-Heisenberg black hole in the presence of PFDM. The precession angles $\Delta\phi_{\text{precession}}$ plotted in this figure are calculated numerically from expression (23) for a given conserved angular momentum value, and the trajectories of orbits are obtained by numerically solving the differential equation (22). The left panels shows the variation of precession angle as the total energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$ changes, while the right panels illustrate the trajectories of particle orbits for a given total energy parameter (It should be noted that here we have substituted $1/2$ in the right hand side of equation (9) such that the energy parameter vanished at infinity, namely $\lim_{r\to\infty}\varepsilon_{\text{energy}}=0$). (a) The upper left panel plots the precession angle of orbits affects by black hole charge $Q$. (b) The upper right panel illustrates the trajectories of orbits for different black hole charge values when total energy equals $\varepsilon_{\text{energy}}=-0.03$ (c) The middle left panel shows the precession angle of orbits affects by dark mater (parameterized by $\lambda_{\text{DM}}$). (d) The middle right panel illustrates the trajectories of orbits for different PFDM parameter values when total energy is $\varepsilon_{\text{energy}}=-0.005$ (e) The lower left panel shows the precession angle of orbits influenced by nonlinear electrodynamics effects / QED effects (parameterized by $a$). (f) The lower right panel illustrates the trajectories of orbits for different nonlinear electrodynamics parameter values when total energy equals $\varepsilon_{\text{energy}}=-0.03$.
The figure 5 presents results on massive object’s bound orbits around the Euler-Heisenberg black hole, including the precession angles in the left panels and the trajectories of orbits in the right panels. The trajectories of massive object’s orbits are obtained by numerically solving the ordinary differential equation (22), and the precession angle of orbits are calculated from expression (23). Correspondingly, the upper, middle and lower panels of figure 5 highlight the influences from black hole electric charge, dark matter (parameterized by PFDM parameter $\lambda_{\text{DM}}$), and nonlinear electrodynamics effects / QED effects (parameterized by nonlinear electrodynamics parameter $a$). In this figure, we present results on the trajectory and precession angle of massive object’s orbits for a selected angular momentum value $J=8M$. The precession angles correspond to other angular momentum values are similar to those for $J=8M$ given in figure 5.
The results in figure 5 show that the black hole electric charge and PFDM can intensely influence the precession angle of massive object’s orbits in the vicinity of Euler-Heisenberg black hole. The upper panels of figure 5 suggests that a larger black hole charge reduces the precession angle of orbits. However, unlike the precession angle, the shape of massive object’s bound orbits are not violently influenced by black hole charge. In the upper right panel of figure 5, the shape and eccentricity of massive object’s orbits calculated for different black hole charge values do not exhibit large differences. On the other hand, the middle panels indicate that the PFDM have great impacts not only on the precession angle of orbits, but also on the shape of massive object’s bound orbits. The middle left panel shows that a negative PFDM parameter can significantly enhance the precession angle. The middle right panel shows that the shape, eccentricity and semimajor-axis length of massive object’s bound orbits (when PFDM parameter takes different values) exhibit notable differences. Moreover, in the middle left panel, the energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$ could not cover the whole region $-0.03\leq\varepsilon_{\text{energy}}\leq 0$ when PFDM parameter $\lambda_{\text{DM}}>-0.4M$. This is due to the local minimum position of effective potential for massive objects. For any massive object’s bound orbits with precession, the total energy $\varepsilon_{\text{energy}}$ in this orbit must be greater than the local minimum of effective potential. The local minimum values of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM (for a selected angular momentum value $J=8.0M$ and several PFDM parameter values) are listed in table 1. Unlike the black hole electric charge and PFDM, the nonlinear electrodynamics effects / QED effects (parameterized by $a$) have very tiny impacts on the trajectories of massive object’s orbits and their procession angles, which can be directly observed from the lower panels of figure 5. This trend is similar to the cases of gravitational deflection and time delay, where the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables.
Table 1: The local minimum position of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM. In this table, the black hole electric charge and non-linear electrodynamics parameter are $Q=0.5M$, $a=0.5M^{2}$, and conserved angular momentum in the massive object’s bound orbit is selected as $J=8M$. In this table, we also substitute $1/2$ in the effective potential given in equation (9) such that the effective potential for massive object vanished at infinity $\lim_{r\to\infty}V_{\text{eff}}(r)=\lim_{r\to\infty}\big{\{}\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}-\frac{1}{2}\big{\}}=0$.
PFDM parameter $\lambda_{\text{DM}}=-0.5M$ $\lambda_{\text{DM}}=-0.4M$ $\lambda_{\text{DM}}=-0.2M$ $\lambda_{\text{DM}}=-0.1M$ $\lambda_{\text{DM}}=0.0M$
local minimum of $V_{\text{min}}=-0.03566$ $V_{\text{min}}=-0.03092$ $V_{\text{min}}=-0.01969$ $V_{\text{min}}=-0.01408$ $V_{\text{min}}=-0.00841$
effective potential
4.4 Black Hole Shadow
This subsection discusses the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM. The critical shadow radius of a black hole can be calculated from the effective potential of photons moving in gravitational field, or through a geometric approach proposed in a recent works [102]. Here, we calculate the radius of unstable photon sphere and the black hole shadow radius based on the geometric approach, in which the photon sphere and its stability are uniquely determined from Gaussian curvature and geodesic curvature in the optical geometry of black hole spacetime. Particularly, for any photon sphere, its geodesic curvature vanishes, which is completely equivalent to find the local extreme point of effective potential [102, 103, 104]
$\kappa_{g}(r=r_{\text{ph}})=0\ \ \Leftrightarrow\ \ \frac{dV_{\text{eff}}(r)}{dr}\bigg{|}_{r=r_{\text{ph}}}=0$ (27)
The unstable photon spheres (or stable photon spheres) are determined by the negative (or positive) sign of Gaussian curvature, which is equivalent to find the local maximum (or local minimum) points of effective potential [102, 103, 104]
$\mathcal{K}(r=r_{\text{unstable}})<0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{unstable}}}<0$ (28a)
$\mathcal{K}(r=r_{\text{stable}})>0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{stable}}}>0$ (28b)
The absolute apparent size of black hole shadow viewed by an observer located at infinity (the asymptotically flat region) is characterized by critical shadow radius. With the unstable photon sphere obtained in equations (27) and (28a), the critical shadow radius of a black hole can be calculated through the relation
$r_{sh}=b_{\text{critical}}=\frac{r_{\text{unstable}}}{\sqrt{f(r=r_{\text{unstable}})}}$ (29)
Figure 6: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels highlight the influences coming from black hole electric charge, dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the critical shadow radius affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$ and $Q=0.5M$). (c) The middle left panel shows the critical shadow radius affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$). (d) The middle right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$ and $Q=0.5M$). (e) The lower left panel presents the critical shadow radius influenced by nonlinear electrodynamics effects (with $\lambda_{\text{DM}}=-0.2M$). (f) The lower right panel shows the black hole shadow in celestial coordinates influenced by black hole electric charge (with $a=0.5M^{2}$ and $\lambda_{\text{DM}}=-0.2M$). In all four panels of this figure, the black hole shadow radius, black hole electric charge, PFDM parameter and nonlinear electrodynamics parameter are given in unit of black hole mass $M$.
The numerical results on black hole shadow for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 6. The left panels plot the critical shadow radius $r_{sh}$ calculated using expression (29), and the right panels illustrate the black hole shadow in celestial coordinates, highlighting the black hole shadow affected by three black hole parameters (the black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$). From figure 6, it is reasonable to draw the following conclusions. Firstly, the increasing of black hole electric charge diminishes the apparent size of black hole shadow. Secondly, the lower left panel shows that the nonlinear electrodynamics parameter $a$ has very tiny influences on the black hole shadow size. For the same electric charge and PFDM parameter, the critical shadow radius calculated with different nonlinear electrodynamics parameter values do not exhibit any obvious differences. Thirdly, when PFDM parameter is not vary large, such as $\lambda_{\text{DM}}\leq 0.5M$ in upper and middle panels of figure 6, the size of black hole shadow is enlarged as PFDM parameter reduced (but this statement is no longer valid when PFDM reaches a large critical point, see figure 7). Particularly, a negative PFDM parameter results in a much larger black hole shadow size.
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) do not produce any naked singularities (see Appendix A for the discussions on horizons of Euler-Heisenberg black hole surrounded by PFDM), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.2M,0.0M,-0.2M,-0.5M$, the critical shadow radius calculated using expression (29) is not continuous at a large black hole charge $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge excesses the critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change, and some branch of unstable photon spheres are disappeared. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is natural to question that the results obtained using expression (29) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
Figure 7: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. This figure exhibits the variation of black hole critical shadow radius with respect to PFDM parameter. (a) The left panel shows the critical shadow radius for Euler-Heisenberg black hole surrounded by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$). (b) The right panel shows the critical shadow radius for Euler-Heisenberg black hole surrounded by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$).
To further explore the dark matter effects on black hole shadow, the figure 7 gives the critical shadow radius changed with respect to PFDM parameter, both in the presence and absence of nonlinear electrodynamics effects. It can be found that when PFDM parameter is smaller than a critical value ($\lambda_{\text{DM}}<\lambda_{\text{critical}}$), the black hole shadow radius reduced with the increasing of PFDM parameter. However, when PFDM parameter excesses this critical value ($\lambda_{\text{DM}}>\lambda_{\text{critical}}$), the black hole shadow radius could be magnified when PFDM parameter increases. This seems to be a universal properties of the PFDM effects on black hole shadows. Some recent studies on charged RN black hole surrounded by PFDM have reported the same conclusion on black hole shadow as we have presented here (the existence of critical point $\lambda_{\text{critical}}$) [50, 49, 51]. Furthermore, form figure 7, the critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values. Since the nonlinear electrodynamics effects have tiny influences on black hole shadow size, the left and right panels of figure 7 exhibit very similar results and tends.
5 Conclusions
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which are obtained from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics and PFDM effects in charge black hole system. It provide us a simply way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects influenced on charged black hole systems. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by black hole parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
Our numerical results presented in this work show that influences from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tend for various gravitational lensing observables (including the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits, and the black hole shadow size). Firstly, the Euler-Heisenberg black hole with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these gravitational lensing observables. Thirdly, the most significant impacts on these gravitational lensing observables coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, the PFDM with a negative dark matter parameter could greatly enhance the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM plays a role of “effective mass”. A negative PFDM parameter contributes to a positive “effective mass” in the spacetime metric, therefore it can enhance the gravitational lensing observables (such as gravitational deflection angle, time delay of light, and precession angle of massive object’s orbits). Moreover, the dark matter medium also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow, when PFDM parameter is smaller than a critical value, the black hole shadow radius reduced as the increasing of PFDM parameter. When PFDM parameter excesses this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
The results in our present work can deepen our understanding of the nonlinear electrodynamics effects / QED effects and dark matter effects on charged black hole systems in the gravitational lensing. Hopefully it may have potential applications for theoretical and observational probing for these dark matter effects and nonlinear electrodynamics effects in black holes and gravitation. Since the dark matter could significantly influence the gravitational lensing observables, the PFDM parameter can be constrained from gravitational lensing observations, by analyzing and comparing the observed data (on gravitational deflection angles, time delays of light, precession angle of bound orbits, apparent black hole shadows) with theoretical predictions, especially with the help of accumulating observational data for supermassive black hole in galaxy centers. However, the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables, so it would be not easy to probe the nonlinear electrodynamics in gravitational lensing observations (unless very high-precision observational data are employed).
Acknowledgment
The authors thanks Song-Lin Lyu and Guang-Zhou Guo for helpful discussion and comments on the numerical schemes. This research was funded by the Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126), the Scientific Research Program of Chongqing Science and Technology Commission (the Chongqing “zhitongche” program for doctors, Grant No. CSTB2022BSXM-JCX0100), the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2020ZDZ027) and the Research and Innovation Team Cultivation Program of Chongqing University of Technology (Grant No. 2023TDZ007).
Abbreviations
The following abbreviations are used in this manuscript:
QED Quantum Electrodynamics
PFDM Perfect Fluid Dark Matter
RN Reissner-Nordström
WIMPs Weakly Interactive Massive Particles
EHT Event Horizon Telescope
Appendix A Horizon of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
This appendix summarizes the results on event horizon of Euler-Heisenberg black hole surrounded by PFDM. The positions of horizon can be solved from the metric component function
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}=0$ (30)
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a=0$ and $\lambda_{\text{DM}}$), the spacetime reduce to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects (or QED effects) and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which are calculated by numerically solving equation (30).
Figure 8: The horizons of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels highlight the influences coming from dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the charged black hole’s horizons affected by PFDM parameter $\lambda_{\text{DM}}$ without the nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel shows the Euler-Heisenberg black hole’s horizons affected by nonlinear electrodynamics effects (parameterized by $a$) without the PFDM. (c) The lower left panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.2M^{2}$). (d) The lower right panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.8M^{2}$). In all four panels of this figure, the horizon radius, black hole electric charge, PFDM parameter and nonlinear electrodynamics parameter are given in unit of black hole mass $M$.
From the upper left panel of figure 8, it can be observed that the PFDM does not affects the structure of horizons in the absence of nonlinear electrodynamics effects (with $a=0$). Without the nonlinear electrodynamics effects / QED effects, the charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. The extremal black hole charge $Q_{\text{extreme}}$ become increased for a smaller PFDM parameter $\lambda_{\text{DM}}$. This can be explained from the “effective mass” role that PFDM played in the spacetime metric (6), as we have explained in section 4. A negative PFDM parameter ($\lambda_{\text{DM}}<0$) could contribute to a positive “effective mass” in the spacetime metric, therefore the corresponding extremal black hole charge can be greater than black hole mass $M$ for a negative PFDM parameter. When the black hole charge excesses the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizon-less and produces a naked singularity in the center point $r=0$.
For Euler-Heisenberg black holes in the presence nonlinear electrodynamics effects (with $a>0$), the structure of horizons can be changed, which are clearly observed in the upper right panel of figure 8. Firstly, the naked singularity no longer exists for large black hole charge when nonlinear electrodynamics is considered. For any given black hole charge, there are at least one horizons near Euler-Heisenberg black hole. Secondly, for a small black hole charge, the spacetime has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel, various curves of horizon radius correspond to different nonlinear electrodynamics parameter values $a$ in the branch $r_{H}\sim 2M$ and $Q<M$ are almost overlapped with each other in the upper right panel, but only the pure RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black dotted curve). Thirdly, for black hole with a relative intermediate electric charge ($0.85M<Q<1.05M$), the Euler-Heisenberg black hole could have three horizons. However, it should be mentioned that the Euler-Heisenberg black hole with a very large nonlinear electrodynamics parameter $a$ (for instance the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons, producing a single horizon for all electric charge values.
When taking into account both nonlinear electrodynamics effects (with $a>0$) and dark matter effects (with $\lambda_{\text{DM}}\neq 0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for any black hole electric charge, which is the same as that for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the wide of three horizon region in the parameter space of black hole charge. The negative PFDM parameter ($\lambda_{\text{DM}}<0$) results in large horizon radius, which can be explained using the “effective mass” role the PFDM played in the spacetime metric. For an intermediate electric charge, a positive PFDM parameter ($\lambda_{\text{DM}}>0$) narrows the three horizon region, while a negative PFDM parameter ($\lambda_{\text{DM}}<0$) widens the three horizon region. For instance, when nonlinear electrodynamics parameter gets value $a=0.2M^{2}$, the PFDM with $\lambda_{\text{DM}}=-0.5M$ produces a much broader three horizon region, compared with those without PFDM plotted in the upper right panel. When $a=0.8M^{2}$, the Euler-Heisenberg black hole in the absence of dark matter does not possess a three horizons region, while the three horizon region emerges for negative PFDM parameter values $\lambda_{\text{DM}}=-0.2M$ and $\lambda_{\text{DM}}=-0.5M$.